# Bayesian Optimization in Language space: An Eval-Efficient AI Self-Improvement Framework

### v1

@@ 1.4 Theoretical Guarantees and Empirical Evaluation
We provide theoretical guarantees on the performance of our approach. We formally show that, under mild regularity conditions, the Best-of-$N$ selection over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2). In other words, T-BoN BO emulates gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model. Since UCB-BO is known to achieve sublinear regret and optimal evaluation-efficiency guarantees, T-BoN BO inherits these theoretical properties up to constant factors and approximation error stemming from textual gradients. In sum, we formally establish that T-BoN BO is theoretically equivalent to the gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret). This equivalence establishes T-BoN BO as a valid, eval-efficient, gradient-based Bayesian optimization procedure in the language space.
In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates. We now demonstrate the empirical performance of our approach using a series of numerical experiments. Concretely, we instantiate T-BoN BO in a digital advertising setting, where the goal is to iteratively refine image-generation prompts for ad creatives so as to maximize ad responsiveness for a population of customers. We test and measure how well and how fast a method optimizes for (i.e., aligns with) the preference distribution induced by LLM-based simulation of personas in the Digital Twins persona dataset called Twin-2k-500 (toubia2025twin). [^3] While LLM-based persona simulation can be biased and may not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution. As such, it can serve as a valid benchmark on which we can compare the performance of different AI-based optimization methods. We demonstrate that T-BoN BO significantly outperforms strong state-of-the-art baselines such as Best-of-N (snell2024scaling) and GEPA (agrawal2025gepa) in terms of evaluation efficiency, i.e., reaching fixed performance with a smaller number of steps. Specifically, we show that:
@@
The rest of the paper is organized as follows. Section 2 discusses the related literature. Section 3 describes the preliminaries related to the ideas discussed in this paper. Section 4 formalizes the AI system and its objective, along with self-improving algorithms. Section 5 introduces the main algorithm proposed in this paper, T-BoN BO, and its components. Section 6 proves that T-BoN BO theoretically emulates the gradient-based UCB Bayesian Optimization. Sections  7 and 8 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines. Finally, Section 9 concludes with a discussion of implications and limitations.
@@ 2 Related Literature
The literature on self-improving AI systems has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI systems must create and learn from their own data. In other words, powerful AI systems “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome). That is, we need to design the self-improving AI system that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs –models, tools, and prompts– in the system to achieve long-term objectives under real-world feedback (wang2025maestro). Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI system has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
In this paper, we study self-improving systems, focusing on tuning prompts, which have recently attracted attention after agrawal2025gepa showed that iterative prompt optimization can be more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek). Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments. Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage. In all those benchmarks, the number of LLM query calls for generating candidate answers was considered as the primary performance bottleneck. Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, i.e., minimizing the number of LLM-generated rollouts. On the other hand, our paper proposes and focuses on the evaluation efficiency objective. That is, instead of focusing on reducing the number of LLM outcome-generation rollouts, we aim to reduce the number of evaluations required to achieve an effective outcome.
@@
Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO). A series of papers focuses on applying LLM to enhance BO for numerical optimization problems. liularge focuses on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization. The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems. Conversely, singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance. In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in LLM-based self-improving AI system design as a Bayesian Optimization problem. In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization. While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection. In contrast to these approaches, our work focuses on self-improving AI systems that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
@@
In terms of method evaluation benchmarks, our paper relates to the methodological literature on ad optimization in digital marketing. A series of papers focused on the efficient execution of ad evaluation, i.e., iteratively improving ad choice efficiently given a fixed set of candidate ads. schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm, which purchased display advertising across 59 websites. geng2020online partitions JD.com’s platform users into disjoint sub-populations and applies a contextual Thompson sampling algorithm to maximize the expected advertiser payoff. aramayo2023multiarmed developed a contextual Thompson sampling algorithm to dynamically determine the list of house ads to display on the homepage of an electronic retailer to maximize the accumulated click-through rate of the ads. ba2022advertising used parametric Thompson Sampling, where a parametric structure links features (media, audience attributes) to conversion likelihood to optimally allocate ad-media and target-audience combinations in high-dimensional settings with low success rates. The main difference between bandit-based papers and ours is the existence of a pre-defined set of candidate arms. Self-improving AI systems focus on creating new solutions by analyzing previous results. On the other hand, bandit methods aim to balance exploration of a predefined set of candidate ads optimally. More broadly, there is also a difference in problem scope. Self-improving AI systems iterate among three steps: generation, evaluation, and analysis. In contrast, bandit methods focus on efficiently executing the evaluation step. Indeed, an AI self-improvement framework’s evaluation step may employ a bandit method to improve efficiency further.
3 Preliminaries
3.1 AI system.
An AI system $\Phi$ can be formally defined as a feed-forward modular workflow, which consists of one or more interconnected Large Language Model (LLM) components. In line with prior work (better_together; khattab2022demonstrate; miprov2; langprobe; agrawal2025gepa), we define an AI system $\Phi$ as a tuple $(M,C,\mathcal{X},\mathcal{Y})$. Here, $M=\langle M_{1},\ldots,M_{|M|}\rangle$ represents the set of LLM-powered components, $C$ is the control logic that orchestrates the workflow and sequences the tasks, and $\mathcal{X}$ and $\mathcal{Y}$ define the system input space and the system output space. Each component $M_{i}=(\pi_{i},\theta_{i},\mathcal{I}_{i},\mathcal{O}_{i})$ is an independent unit where $\pi_{i}$ is the prompt, $\theta_{i}$ represents the fixed weights of the underlying language model, and $\mathcal{I}_{i},\mathcal{O}_{i}$ specify the component’s input space and output space.
We denote the prompts in the whole system $\Pi=\left\langle\pi_{1},\ldots,\pi_{|M|}\right\rangle$. As the prompts $\Pi$ are what we system designers can control for the system $\Phi$, we denote $\Phi$ as the functional $\Phi(\cdot)$ when we would like to emphasize system $\Phi$’s dependency on prompts $\Pi$. Accordingly, we call $\Phi(\Pi)$ the ‘system outcome’ corresponding to prompt $\Pi$. Note that $\Phi(\Pi)$ is itself a function that maps $\mathcal{X}$ to $\mathcal{Y}$, i.e., $\Phi(\Pi):\mathcal{X}\mapsto\mathcal{Y}$.
Figure 2 illustrates an example AI system with five LLM-powered components. Here, its control logic $C=\{\mathcal{X}\rightarrow\left\{M_{1},M_{2}\right\},M_{1}\rightarrow M_{3},M_{2}\rightarrow M_{4},\left\{M_{3},M_{4}\right\}\rightarrow M_{5}\rightarrow\mathcal{Y}\}$ describes that the the system input $\mathcal{X}$ becomes the component input for $M_{1}$ and $M_{2}$, $M_{3}$’s input is $M_{1}$’s output, $M_{4}$’s input is $M_{2}$’s output, $M_{5}$’s input is $M_{3}$’s output and $M_{4}$’s output, and the final output $\mathcal{Y}$ is the output of $M_{5}$. The system $\Phi$’s input is the prompt $\Pi=\langle\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\rangle$.
[…]
Figure 2: An example AI system with control logic $C=\{\mathcal{X}\rightarrow\left\{M_{1},M_{2}\right\},M_{1}\rightarrow M_{3},M_{2}\rightarrow M_{4},\left\{M_{3},M_{4}\right\}\rightarrow M_{5}\rightarrow\mathcal{Y}\}$ and prompt $\Pi=\langle\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\rangle$.
3.2 TextGrad: Textual gradients in language space.
A central question for extending gradient-based Bayesian optimization to text is how to define a usable notion of “gradient” in language. A recent influential paper by (textgrad) addresses this problem by proposing a framework called the TextGrad; and it has has attracted broad interest across multiple research communities. Instead of computing numerical derivatives, TextGrad asks a critic LLM to propose targeted edits that would improve a candidate text. These edits act as local, directional signals, i.e., textual gradients, that guide search in a discrete language space. Importantly, this approach does not require embedding relaxations or external surrogate models, making it directly applicable to prompt optimization and other text-centric tasks.
For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal.
Empirical studies provided in textgrad show that textual gradients enable consistent improvements across diverse scientific domains. In programming tasks, they raise success rates on challenging code-completion benchmarks. In knowledge-intensive reasoning, they improve accuracy on graduate-level QA datasets. In applied domains, they have been used to optimize drug-like molecules (via text encodings) and to refine radiotherapy planning prompts for clinical objectives. These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in and over language, justifying our use of gradient-style search in the discrete text space.
3.3 Bayesian Optimization.
Bayesian Optimization (BO) is a framework for maximizing an unknown and expensive-to-evaluate objective function under a limited evaluation budget. Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts. BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$. For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
The next evaluation point is selected by maximizing an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$. Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty). A widely used choice is the Upper Confidence Bound (UCB) acquisition function is:
$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (1)
where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information; wilson2018maximizing). In many applications, the maximizer of $A_{\beta}$ is not available in closed form, and one instead employs gradient-based BO, which performs local ascent on $A_{\beta}$ (often from multiple random initializations) to obtain approximate maximizers (wilson2018maximizing; papenmeier2025understanding). In gradient-based BO, each evaluation is followed by multiple gradient steps that identify local optima of the acquisition function. Gradient-based BO methods have shown surprising effectiveness for high-dimensional problems under random multi-starts and careful choice of gradients (papenmeier2025understanding).
In our setting, we do not construct an explicit surrogate model; rather, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs realized within the LLM and design our method to emulate the behavior of gradient-based UCB in this implicit space.
4 Overview of Self-Improving AI Algorithms
@@
Consider a system designer’s decision making problem, where she chooses the prompt set $\Pi$ for the LLM-based AI system $\Phi$. [^4] We call $\Pi$ in a plural form ‘prompts’ or ‘prompt set’ instead of a singular form ‘prompt’, as there are often multiple prompt components in one system. See AI system paragraph in Section 3.1 for a detailed description of an AI system. [^5] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek). In this paper, we focus on prompt optimization, which fixes those internal parameters throughout. We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt set $\Pi$. In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt set $\Pi$. As a system, $\Phi(\Pi)$ has inputs and corresponding outputs. Specifically, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ to a distribution over an output space $\mathcal{Y}$. For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued vector output of the system at input $x$. In a digital marketing setting, input $x\in\mathcal{X}$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$.
@@
4.2 Critic LLM and Self-improving AI algorithm
Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi)$ through iteratively analyzing the previous history, proposing new prompts, and observing their scores. This is often enabled by assuming an external critic LLM, denoted $M_{\text{critic}}$, that can analyze previous data and propose the next solution.
Critic LLM.
Specifically, a critic LLM is assumed to have the following capabilities:
  • 1)
    The critic LLM can look at $\Phi(\Pi_{1})$ and $\Phi(\Pi_{2})$ side by side and predict which one will be better.
  • 2)
    The critic LLM can read the history of past attempts and summarize the main lessons as a short list of rules or guidelines.
  • 3)
    The critic LLM can suggest a small set of prompt edits that are likely to improve the current prompt.
In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), and (3) propose prompt improvements based on those guidelines (textgrad; li2025test).
Self-improving AI algorithm
Given an initial prompt defined as $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$. At the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores and then selects a new prompt set $\Pi_{t}$, deploys the corresponding system $\Phi(\Pi_{t})$, and receives feedback about its performance. At each algorithm iteration $t$, we define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,

### v2

@@ 1.5 Theoretical Guarantees and Empirical Evaluation
We provide theoretical guarantees on the performance of our approach. We formally show that, under mild regularity conditions, the Best-of-$N$ gradient over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2). In other words, T-BoN BO emulates parallel gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model. Since UCB-BO is known to achieve sublinear regret and optimal evaluation-efficiency guarantees, T-BoN BO inherits these theoretical properties up to constant factors and approximation error stemming from textual gradients. In sum, we formally establish that T-BoN BO is theoretically equivalent to the parallel gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret). This equivalence establishes T-BoN BO as a valid, eval-efficient, parallel gradient-based Bayesian optimization procedure in the language space.
In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since it does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates. We now demonstrate the empirical performance of our approach using a series of numerical experiments. Concretely, we instantiate T-BoN BO in a digital advertising setting, where the goal is to iteratively refine image-generation prompts for ad creatives so as to maximize ad responsiveness for a population of customers. We test and measure how well and how fast a method optimizes for (i.e., aligns with) the preference distribution induced by LLM-based simulation of personas in the Digital Twins persona dataset called Twin-2k-500 (toubia2025twin). [^2] While LLM-based persona simulation can be biased and may not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution. As such, it can serve as a valid benchmark on which we can compare the performance of different AI-based optimization methods. We demonstrate that T-BoN BO significantly outperforms strong state-of-the-art baselines such as Best-of-N (snell2024scaling) and GEPA (agrawal2025gepa) in terms of evaluation efficiency, i.e., reaching fixed performance with a smaller number of steps. Specifically, we show that:
@@
The rest of the paper is organized as follows. Section 2 discusses the related literature. Section 3 formalizes the AI system and its objective, along with self-improving algorithms. Section  4 describes the concept of Best-of-N gradient and proves that it theoretically emulates the gradient of the UCB acquisition function. Section 5 introduces the main self-improving AI algorithm proposed in this paper, T-BoN BO. Sections 6 and 7 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines. Finally, Section 8 concludes with a discussion of implications and limitations.
@@ 2 Related Literature
The literature on self-improving AI has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI must create and learn from its own data. In other words, powerful AI “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome). That is, we need to design the self-improving AI that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs –models, tools, and prompts– in the system to achieve long-term objectives under real-world feedback (wang2025maestro). Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro). In this paper, we focus on studying self-improving systems based on tuning prompts.
@@
Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO). A series of papers focuses on applying LLM to enhance BO for numerical optimization problems. liularge and cisse2025language focus on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization. The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems. Conversely, singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance. In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in formulating LLM-based self-improving AI system design as a Bayesian Optimization problem. In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization. While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection. In contrast to these approaches, our work focuses on self-improving AI that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
@@
In terms of method evaluation benchmarks, our paper relates to the methodological literature on ad optimization in digital marketing. A series of papers focused on the efficient execution of ad evaluation, i.e., iteratively improving ad choice efficiently given a fixed set of candidate ads. schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm, which purchased display advertising across 59 websites. geng2020online partitions JD.com’s platform users into disjoint sub-populations and applies a contextual Thompson sampling algorithm to maximize the expected advertiser payoff. aramayo2023multiarmed developed a contextual Thompson sampling algorithm to dynamically determine the list of house ads to display on the homepage of an electronic retailer to maximize the accumulated click-through rate of the ads. ba2022advertising used parametric Thompson Sampling, where a parametric structure links features (media, audience attributes) to conversion likelihood to optimally allocate ad-media and target-audience combinations in high-dimensional settings with low success rates. The main difference between bandit-based papers and ours is the existence of a pre-defined set of candidate arms. Self-improving AI focuses on creating new solutions by analyzing previous results. On the other hand, bandit methods aim to balance exploration of a predefined set of candidate ads optimally. More broadly, there is also a difference in problem scope. Self-improving AI iterates among three steps: generation, evaluation, and analysis. In contrast, bandit methods focus on efficiently executing the evaluation step. Indeed, an AI self-improvement framework’s evaluation step may employ a bandit method to improve efficiency further.
3 Setup and Background
@@
Consider a system designer’s decision-making problem, where she chooses a prompt $\Pi$ for the LLM-based AI system $\Phi$. [^3] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek). In this paper, we focus on prompt optimization, which fixes those internal parameters throughout. We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt $\Pi$. In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt $\Pi$. As a system, $\Phi(\Pi)$ has inputs and corresponding outputs. Specifically, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ to a distribution over an output space $\mathcal{Y}$. For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued vector output of the system at input $x$. In a digital marketing setting, input $x\in\mathcal{X}$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$.
@@
3.2 Self-improving AI algorithm
Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi):=\mathbb{E}[r(y(\Phi(\Pi),x))]$ through iteratively analyzing the previous history, proposing a new prompt, and observing its noisy score. Formally, starting from the initial prompt $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$; at the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores, then selects a new prompt $\Pi_{t}$, deploys the corresponding system outcome $\Phi(\Pi_{t})$, and receives a score feedback on its performance. We define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,